# Space as Time Through Neuron Position Learning

### v1

@@ 2.4 Unconstrained position learning converges on more local solutions
Figure 3: A pruning effects on classification accuracy on SHD. Pruning is less harmful for the position learning network. B delay distributions post training. For the non-spatial network, most delays are still zero after training, but we also end up with a few very long delays that are non visible since they are significantly outnumbered by the zero delays. For the spatial network delays become spread out in the lower range. C delay length based pruning effects on modularity. Position learning network achieves higher modularity. D delay length-based pruning effects on small-worldness. Position learning network achieves higher small-worldness.
@@
We also study networks with no regularisation, which, of course, is not biologically realistic, as connecting to nearby neurons in the brain is always ‘cheaper’. We were interested, however, in what solutions the network finds under no constraints, and if we still converge on solutions where certain synapses are more important than others. We found that ‘classic’ weight magnitude-based pruning and random pruning had very similar effects on both non-spatially embedded networks and those where the positions were learnt. Taking inspiration from biology and emphasising short connections, we pruned away the top $x\%$ (ranging from $5\%$ to $30\%$) longest connections. The learnt delay distributions of non-spatial and spatial networks seem to also hint (see Figures 3), that our pruning studies show that spatial networks converge on solutions that are more reliant on short local connections (see Figure 3). Given that this was not enforced through any explicit regularisation, this is a surprising finding, particularly since long delays tend to be beneficial for temporal tasks (Sun et al., 2025). These findings not only align with neuroscience properties but they have implementation benefits as well. In digital neuromorphic systems, delays are usually implemented through costly dynamic memory buffers, so shortening delays can be beneficial. Furthermore, most neuromorphic chips also have an upper bound on delays (Davies et al., 2021). Given that random and magnitude-based pruning did not show such differences, seemingly spatial networks are not simply more robust but are more ‘local’. We also tested robustness through neuron ablation studies, and adding noise to delays as well and we did not observe a significant difference between spatial and non-spatial architectures.
@@
Since delay length based pruning has less impact on the performance of spatial architectures, we were curious to see how the network topology changes after such pruning. We again study modularity in our networks, and also small-worldness. Note, that neither of them take physical distances between nodes into account. Figure  3 shows the effect of delay length pruning on modularity and small-worldness. As a baseline, we calculated the modularity of a random matrix, and found that spatial architectures show a clear increase in modularity with pruning, while non-spatial architectures show a similar trend as pruning a random matrix. Similarly, small-worldness increases more rapidly in spatial networks (with the same connectivity ratio). We observe the higher accuracy, modularity and small-worldness after pruning in spatial architectures, without explicitly enforcing any of these structures to emerge.
@@ 3 Discussion
Here we introduced a method to study the intertwinement of space and time in neural networks. While neuron position learning has been explored before, prior studies have mainly emphasized parameter efficiency (Erb et al., 2025) or performance improvements (Vasilache et al., 2025) on machine learning benchmarks. Here, we instead focus on how the structure or learnt networks changes, when space and time are intrinsically connected.
We find that neuron positions in the hidden layer correlate with input-layer connectivity patterns. However, in our simple classification setting, it remains unclear what functional roles clusters of nearby neurons serve. Long delays between modules specialized for different subtasks may be advantageous when temporal alignment across these modules is required for solving a more complex task. Extending this analysis to modular tasks (Béna & Goodman, 2025) with an added temporal structure, to multimodal settings, or multitask problems (Vafidis et al., 2024) presents an interesting avenue for future work.
In our examples we focus on the inputs’ representation in the hidden layer. Of course it would be interesting to see if neurons (or even groups of neurons) specialise on particular inputs, but since we are compressing from the input to the hidden layer (700 input neurons to 128 neurons), we would not expect specialisation like that to emerge. As mentioned before, we study small networks, to make sure that delays become useful for for the task. We could also of course constrain the networks in other ways such as by constraining neurons to only spike once. Another solution would be to turn to tasks where we have a small input population with complex information, such that recurrent delays are still crucial even when we expand from the input to the hidden layer.
@@
Achterberg et al. (2023) investigated networks with fixed neuron positions. If positions were instead allowed to be optimized, all neurons would likely converge toward the centre. In such a case, the benefits of modularity and small-world structure would vanish, since distance acts only as a cost. Once proximity carries no penalty, the network is free to adopt any connectivity pattern needed to solve the task. In contrast, by introducing a benefit of distance, our framework prevents this collapse and encourages spatially distributed structures. Similarly, in biological neural networks, most connections are short-range, yet the overall average connection length remains higher than the physical minimum (Bassett et al., 2010). During development, neuronal migration (de Graaf-Peters & Hadders-Algra, 2006) contributes to functional circuit formation. However, biological constraints such as each neuron occupying a physical volume (volume exclusion) prevents neurons from being too close to one another. In our model, neurons can theoretically be arbitrarily close, provided they are not coincident. Introducing an explicit repulsive force between two neurons would be straightforward to formalize but challenging to extend to more complex circuits—offering another interesting direction for future work. Furthermore, now we even have studies that track neurons position changes over time (Akarca et al., 2025), allowing for direct comparisons between the presented algorithm and empirical data.
As we have shown, with position learning, networks tend to converge to more local solutions, where the core computations occur within small-world and modular topologies. While these findings are interesting from a neuroscience perspective, such solutions may also be beneficial for neuromorphic implementations. For example, the chip introduced by Dalgaty et al. (2024) relies on small-world connectivity. Weber et al. (2025) demonstrated that the structural plasticity algorithm DEEP R (Bellec et al., 2017) can be extended to account for system constraints (e.g., small-world topology). Inspired by this, introducing distance awareness with position learning to DEEP R could yield interesting network geometries, as neuron positions would no longer need to be constrained to a hypersphere (as we did here) to comply with system constraints.
Keller et al. (2024) outlines the importance of spacetime neural representation both for theoretical neuroscience and improved generalisation and long-term working memory in AI. Travelling waves have been shown to facilitate global information integration (Jacobs et al., 2025). Izhikevich & Hoppensteadt (2009) even proposed to implement such methods without synapses (or in other other words, fully connected networks with uniform synapses), showing that such set up can perform temporal pattern recognition, reverberating memory, temporal signal analysis and basic logical operations. Because distance-dependent delays in large spiking networks naturally give rise to such waves (Davis et al., 2021), investigating position learning in this context is a particularly promising direction.
Appendix A Methods
@@
For this work, the original EventProp backward dynamics (Wunderlich & Pehle, 2021; Nowotny et al., 2025) do not change. The weight update rule with delays (Mészáros et al., 2025) is defined as
@@
where $\Lambda_{1}$ is the regularisation strength. If we introduce the distance cost as well, the equation is
$\frac{{\rm d}\mathcal{L}}{{\rm d}w_{ji}}=-\tau_{s}\sum_{t_{k}fromi}\lambda_{I,j}\Big|_{t_{k}+d_{ji}}+\Lambda_{1}sign(w_{ji})d_{ji}.$ (9)
@@
We define small-worldness as
$\sigma=\frac{c/c_{rand}}{l/l_{rand}},$ (11)
where $c$ and $c_{rand}$ are clustering coefficients, and $l$ and $l_{rand}$ are the characteristic path lengths of the tested network and the mean of $1000$ random networks with the same size and density. In the fully connected network $\sigma\approx 1$.
The Shannon entropy of the weight matrix $W$ is defined as
@@
We define the communicability matrix as
@@
and we can calculate the communicability entropy with the combination of the previous two equations,
@@
For this work we used the mlGeNN library (Knight & Nowotny, 2023). We implemented support for weight regularisation, the position gradient update function, and axonal delays.
Appendix B Architecture comparisons
Figure 4: Architecture comparisons

### v2

@@ 2.5 Unconstrained position learning converges on more local solutions
Figure 6: Results of pruning longest connections. We depict the mean of the results over 20 random seeds, error bars represent standard deviation. (A) Pruning effects on classification accuracy on SHD. Pruning is less harmful for the position learning network. (B) Delay distributions post training. For the non-spatial network, most delays are still zero after training, but we also end up with a few very long delays that are not visible since they are significantly outnumbered by the zero delays. For the spatial network delays become spread out in the lower range. (C) Delay length based pruning effects on modularity. Position learning network achieves higher modularity. (D) delay length-based pruning effects on clustering. Position learning network achieves higher clustering.
@@
Finally, we also study networks with no regularisation. This is, of course, not biologically realistic as connecting neurons in the brain always has a ‘cost’ (particularly if they are distant). However, we were interested in what solutions the network finds without constraints and, whether networks still converge on solutions where certain synapses are more ‘important’ than others. We found that ‘classic’ weight magnitude-based pruning and random pruning had very similar effects on both non-spatially embedded networks and those where positions were learnt. Taking inspiration from biology and emphasising short connections, we pruned away the top $x\%$ (ranging from $5\%$ to $30\%$) longest connections. The learnt delay distributions of non-spatial and spatial networks shown in figure 6A seem to hint that spatial networks converge on solutions which are more reliant on short local connections. Given that this was not enforced through any explicit regularisation, this is a surprising finding, particularly since long delays tend to be beneficial for temporal tasks (Sun et al., 2025). These findings not only align with findings from neuroscience (Swadlow, 1985) but also have implementation benefits, as, in digital neuromorphic systems, delays are usually implemented through costly dynamic memory buffers whose size has a strict upper bound (Davies et al., 2021). Therefore, shortening delays can be beneficial. Given that random and magnitude-based pruning did not show such differences, seemingly spatial networks are not simply more robust but are more ‘local’. We also tested robustness through neuron ablation studies and adding noise to delays, neither of which resulted in a significant difference between spatial and non-spatial architectures.
@@
Since delay-based pruning has less impact on the performance of spatial architectures, we were curious to see how the network topology changes after pruning. Therefore, we recalculate modularity and clustering in our pruned networks (since the connectivity is now sparse, we disregard the weights and measure these metrics on the binary connectivity matrices using equations (12) and (13) in the Appendix). Figure  6C-D shows the effect of delay-based pruning on modularity and clustering. As a baseline, we calculated the modularity of a random matrix, and found that spatial architectures show a clear increase in modularity with pruning, while non-spatial architectures show a similar trend as pruning a random matrix. Similarly, clustering increases more rapidly in spatial networks (with the same connectivity ratio). After pruning, we observe higher accuracy, modularity and clustering in spatial architectures, without explicitly enforcing any of these structures to emerge.
@@ 3 Discussion
Here, we have introduced a method for studying the intertwinement of space and time in spiking neural networks. While neuron position learning has been explored before, prior studies have mainly emphasised parameter efficiency (Erb et al., 2026; Tohouri, 2025; Landsmeer et al., 2025) or performance improvements (Vasilache et al., 2025) on machine learning benchmarks. Here, we instead focus on how the structure of learnt networks changes when space and time are intrinsically connected through delays.
@@
Achterberg et al. (2023) investigated networks with fixed neuron positions. If positions were instead allowed to be optimised, since distance acts only as a cost, all neurons would likely converge toward the centre, meaning the benefits of modularity and clustered structure would vanish. Once proximity carries no penalty, the network is free to adopt any connectivity pattern needed to solve the task. In contrast, by introducing a benefit of distance, our framework prevents this collapse and encourages spatially distributed structures. Similarly, in biological neural networks, most connections are short-range, yet the overall average connection length remains higher than the physical minimum (Bassett et al., 2010). During development, neuronal migration (de Graaf-Peters and Hadders-Algra, 2006) contributes to functional circuit formation. However, biological constraints such as each neuron occupying a physical volume (volume exclusion) prevents neurons from being too close to one another. In our model, neurons can theoretically be arbitrarily close, provided they are not coincident. Introducing an explicit repulsive force between two neurons would be straightforward to formalise but challenging to extend to more complex circuits – offering another interesting direction for future work. Furthermore, now we even have studies that track neurons’ position changes over time (Akarca et al., 2025), allowing for direct comparisons between the presented algorithm and empirical data.
We find that the positions of neurons in the hidden layer correlate with input to hidden connectivity patterns. From a mechanistic interpretability perspective, these correlations point to concrete computational mechanisms. In our simple classification setting, the precise functional roles of clusters of nearby neurons remain unclear. However, this is likely due to the limited temporal structure of the task rather than a lack of internal specialisation. Long delays between modules specialised for different subtasks may be advantageous when temporal alignment across these modules is required for solving a more complex task. Therefore, extending this analysis to modular tasks (Béna and Goodman, 2025) with an added temporal structure, to multimodal settings, or multitask problems (Vafidis et al., 2025) provide a promising avenue for extending this mechanistic analysis. In our examples, we focus on the representation of inputs in the hidden layer. Of course, it would also be interesting to see if neurons (or even groups of neurons) specialise on particular inputs, but since we are compressing from the input to the hidden layer (700 input neurons to 128 hidden neurons), we would not expect specialisation like that to emerge. As mentioned before, we study small networks to make sure that delays become useful for the task. We could also constrain the networks by only allowing hidden neurons to spike once or by turning to tasks with low-dimensional but complex inputs, meaning that recurrent delays remain crucial even with a larger hidden layer.
Depending on the constraints imposed during training, our networks consistently converge to structured spatial organisations which have distinct practical implications for neuromorphic hardware. Networks trained with position learning converge to solutions that exploit synaptic delays while keeping them short, which is desirable for digital neuromorphic implementations (Davies et al., 2021). Imposing L1 regularisation scaled by interneuron distance leads to highly modular architectures. Such modular designs are particularly well suited to resource-constrained autonomous agents, which naturally favour modular organisation (Bartolozzi et al., 2022). Interestingly, we observe that while distance cost leads to clustered architectures, imposing pure L1 regularisation achieves even higher values. This is directly relevant for recent neuromorphic hardware designs, such as Mosaic (Dalgaty et al., 2024), which explicitly relies on clustered connectivity. Weber et al. (2025) demonstrated that the structural plasticity algorithm DEEP R (Bellec et al., 2018) can be extended to account for system constraints (e.g., small-world topology). Inspired by this, introducing distance awareness with position learning to DEEP R could yield interesting network geometries, as neuron positions would no longer need to be constrained to a hypersphere, but merely comply with system constraints. Furthermore, our results seem to imply that using distance-dependent delays, we could find architectures that are inherently more clustered, and thus could make the best use of hardware which requires clustered connectivity.
In this work, we deliberately focus on small networks in order to enable detailed analysis and clear attribution of observed effects. However, this choice does not reflect an inherent limitation of SNNs, which have been successfully trained at large-scales (Zhu et al., 2024). Scaling is particularly interesting in the present context, as spatially local connectivity and distance-dependent delays are key ingredients for the formation of travelling waves (Davis et al., 2021). Keller et al. (2024) outlines the importance of spacetime neural representation both for theoretical neuroscience and improved generalisation and long-term working memory in AI. Travelling waves, in particular, have been shown to facilitate global information integration (Jacobs et al., 2025; Keller and Welling, 2023; Liboni et al., 2025).
Software and Data
[…]
Acknowledgments
This work was funded by the EPSRC (grants EP/V052241/1 and EP/S030964/1) and the EU (grant no. 945539). Additionally, BM was funded by a Leverhulme Trust studentship and by The Alan Turing Institute, and DA was funded by an Imperial College Research Fellowship, Schmidt Sciences Fellowship and Templeton World Charity Foundation, Inc (funder DOI 501100011730) under the grant TWCF-2022-30510. Compute time was provided through Gauss Centre for Supercomputing (application numbers 21018, 30182 and 61883) and EPSRC (grant number EP/T022205/1) and local GPU hardware was provided by an NVIDIA hardware grant award.
Impact Statement
This paper presents work aiming to advance the field of Machine Learning. If developed in further applied research, there may be many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Comparing accuracy of networks
To ascertain that position learning is useful, we first compare position learning networks of hidden layer sizes between 32 and 128 and embeddings of 2, 3 and 4 dimensions, against networks with non-spatially constrained synaptic, and axonal delays, and networks without any delays. Axonal delays mean that all connections originating from the same neuron have an identical delay, i.e. the delay matrix becomes a delay vector. We could have additionally compared against networks with fixed random delays, but it raises the question of how to initialise these delays. In recurrent connections, this is a particularly complex and understudied question. We illustrate the comparisons of the networks in figure 7.
Appendix B Preferred positions
We calculate the preferred position of each input neuron $i$ in the $x$ dimension with the following equation:
${x^{\mathrm{pref}}_{i}}=\frac{\sum_{j}{w_{ij}}\,{x_{j}}}{\sum_{j}{w_{ij}}}$ (7)
where $w_{ij}$ is the learnt synaptic weight between input neuron $i$ and hidden neuron $j$. Instead of using the learn weights, we can also calculate ‘sensitivities’ ($\tilde{w}_{ij}(t)$) , derived from neuron activities:
$\tilde{w}_{ij}(t)=\bf{1}_{\{t={t_{k}^{i}}\;\land\;|t-{t_{n}^{j}}|<\Delta\}},$} (8)
where $t_{k}^{i}$ denotes the time of the k-th spike of input neuron $i$, and $t_{n}^{j}$ denotes the time of the n-th spike of hidden neuron $j$.
Figure 7: Network comparisons. We depict the mean of the results over 20 random seeds, error bars represent standard deviation.
With a fixed number of trainable parameters, we observe that networks with delays always outperform networks without them. As network size (i.e. number of hidden neurons) increases, the performance gaps decrease. This is not surprising, as all architectures have recurrent connections, which, similar to delays, can serve as a temporal memory. As the network gets more and more neurons to store information in, delays become less and less useful. Spatially embedded networks perform similarly to networks with unconstrained delay learning. Axonal delays, however, perform at least as well as the synaptic delays derived from the Euclidean distances. Note, that the error bars are large because we study small networks, to make sure that there is a clear benefit of delays.
Appendix C Regularisation
@@
For this work, the original EventProp backward dynamics (Wunderlich and Pehle, 2021; Nowotny et al., 2025) do not change. The weight update rule with delays (Mészáros et al., 2025) is defined as
@@
where $\Lambda_{1}$ is the regularisation strength. If we introduce the mean scaled distance cost as well, the equation is
$\frac{{\rm d}\mathcal{L}}{{\rm d}w_{ji}}=-\tau_{s}\sum_{t_{k}fromi}\lambda_{I,j}\Big|_{t_{k}+d_{ji}}+\Lambda_{1}sign(w_{ji})d_{ji}/\bar{D},$ (11)
where $\bar{D}$ is the mean of the delay matrix. The regularisation term was omitted from the neuron position updates (i.e. the gradient of the regularisation term was directly added to weight gradients and not to the loss function), so that only the topology of the network was effected, and not the geometry.
Figure 8: We depict the mean of results over 20 random seeds, error bars represent standard deviation. (A) Classification accuracy over increasing regularisation strength. (B) Connection sparsity over increasing regularisation strength.
Appendix D Modularity
@@
Appendix E Small-worldness
We define clustering as
$C=\frac{1}{N}\sum c_{i,j},$ (13)
where $c_{i,j}$ is the number of closed triangular motifs including node $i$. We normalise this by the clustering of a null-model of the same density:
$\Gamma=\frac{C}{C_{rand}}.$ (14)
The weighted clustering is defined as:
$C_{w}=\frac{1}{k_{i}(k_{i}-1)}\sum_{j,k}(w_{ij}w_{jk}w_{ik})^{1/3},$ (15)
normalised by a null-model which is constructed by random permutations of the weight matrix. Figure 4 shows increased weighted clustering with increased regularisation stength, and figure 6 shows increased clustering with pruning the spatial network. In the binary case, we define the path length as:
$L=\frac{1}{N}\sum l_{i,j},$ (16)
where $l_{i,j}$ is the shortest path between nodes $i$ and $j$, and again, we normalise with a null-model of the same density:
$\Lambda=\frac{L}{L_{rand}}.$ (17)
The definition of path length in the weighted case is:
$L_{w}=\frac{1}{N(N-1)}\sum_{i\neq j}\frac{1}{w_{i,j}},$ (18)
normalised again by the average path-length of random permutations of the weight matrix. With these two equations, we can define small-worldness as:
$\sigma=\frac{\Gamma}{\Lambda}.$ (19)
By definition, we consider a network small-world if $\sigma>1,\Gamma>1$ and $\Lambda\approx 1$. We observed high values of $\Gamma$ in our experiments, thus if $\Lambda\approx 1$, we have small-world networks.
Figure 9: Path length after pruning longest delays. We depict the mean of the results over 20 random seeds, error bars represent standard deviation.
We observe in figure 9 that for the non-regularised, pruned networks the path length insignificantly decreases from the initial value of $1$, thus, we conclude pruning the spatial network yields leads to small-world topologies.
Figure 10: Weighted path length with various regularisation strengths. We depict the mean of the results over 20 random seeds, error bars represent standard deviation.
In contrast, for the regularised networks (figure 10) we can see that the weighted path length significantly decreases with stronger regularisation. This might be due to the network converging on solutions with multiple smaller populations of neurons operating separately in the hidden layer. Thus, these networks are not small-world.
Appendix F Wiring efficiency
We define the wiring efficiency by
$\text{wire}=\langle W,D\rangle_{F},$ (20)
and normalise these values between 0 and $1$, by calculating the maximum and minimum possible wiring, with the learnt network geometry. By sorting the weights in an ascending order, and the delays in a descending order and again calculating the product, we get the minimum possible wiring ($\text{wire}_{\text{min}}$), and by ordering both in a descending (or ascending) order we get the maximum possible wiring ($\text{wire}_{\text{max}}$.We then normalise the efficiency with the following equation:
$\text{wire}_{\text{norm}}=1-\frac{\text{wire}-\text{wire}_{\text{min}}}{\text{wire}_{\text{max}}-\text{wire}_{\text{min}}}$ (21)
Appendix G Entropy
Shannon entropy quantifies the amount of unpredictability. A highly predictable system has low entropy and will exhibit a degree of clustering around particular weight values. An unpredictable system has high entropy and will be more uniform in its distribution. While the previously introduced modularity metric Q shows the extent to which the network can be partitioned into subcommunities, it does not measure the degree of uncertainty of the weight distribution. The Shannon entropy of the weight matrix $W$ is defined as
@@ Appendix G Entropy
In Figure 11 we observe, that position learning networks with distance cost achieve the lowest entropy, in contrast to others which increase in entropy when regularisation is strong.
Figure 11: Shannon entropy over various regularisation strengths. We depict the mean of the results over 20 random seeds, error bars represent standard deviation.
In this work, we are interested in the effect of space on network structure, and thus only introduced the distance cost to the regularisation term. Achterberg et al. (2023) and Sheeran et al. (2024) also introduced a communicability cost, which we did not employ here. However, we can still measure the communicability entropy, similarly to Sheeran et al. (2024). We define the communicability matrix as
@@
and we can calculate the communicability entropy as
@@
Figure 12: Communicability entropy over various regularisation strengths. We depict the mean of the results over 20 random seeds, error bars represent standard deviation.
In figure 12, we can observe that networks with the introduced distance cost yield a lower entropy. In contrast to the Shannon entropy, the spatial network without the distance cost shows a decrease in communicability entropy as the regularisation strength increases. These architectures yield a small number of highly communicable connections, with an increasingly ordered topology.
Appendix H Implementation
For training our models we used the mlGeNN library (Knight and Nowotny, 2023). We implemented support for weight regularisation, the position gradient update function, and axonal delays. For our graph measures we used the Brain Connectivity Toolbox (Rubinov and Sporns, 2010).